OpenStudy (anonymous):

Half of Mr.Mcnair's age minus twice his daughter's age is 17. Double Mr.Mcnair's age plus his daughter's age is 104. How old is Mr.mcnair? how old is his daughter?

zepdrix (zepdrix):

Half of Mr's age(1/2M), minus(-), twice his daughters age(2D), is(=), 17. (1/2)M - 2D = 17 Double Mr's age(2M), plus(+), his daughter's age(D), is(=), 104. 2M + D = 104 With these types of problems, it's important that you are able to recognize little keywords. For example IS means EQUALS. :) So these are our two equations we'll be starting with: (1/2)M - 2D = 17 2M + D = 104

zepdrix (zepdrix):

Try to solve them similar to the way we did the last one, it should work out nicely. :D Lemme know what you get. (Or if you get stuck)

OpenStudy (anonymous):

thanks, and now we do substitution like the other one?

zepdrix (zepdrix):

Like the last one, we're going to get rid of that 1/2 first. Then we'll try to Add or Subtract the 2 equations :D

OpenStudy (anonymous):

so we multiply the first equation with a 17?

zepdrix (zepdrix):

No, we want to multiply by a number that will get rid of the 1/2. 2 is a good number for that. In the last problem, we had to deal with 1/2 and 1/3. So we needed a number that would get rid of BOTH fractions. 6 was lowest number that could deal with both :)

OpenStudy (anonymous):

oh so do we multiply the first one only or the second one too?

zepdrix (zepdrix):

Mmm let's just multiply the first equation by 2 :) the second one seems fine.

zepdrix (zepdrix):

What do you get for equation 1 after you do that? c:

OpenStudy (anonymous):

1m-4d=34

zepdrix (zepdrix):

yay \c:/

OpenStudy (anonymous):

and now do we multiply to get rid of the variables?

zepdrix (zepdrix):